OpenStudy (anonymous):

How do I solve this absolute value equation for x: |x-5| + |2-2x| = 11

OpenStudy (blockcolder):

|x-5|=11-|2-2x| \[11-|2-2x|= \begin{cases} x-5, &\text{if }x\geq5\\ 5-x &\text{if }x<5 \end{cases}\] Now separately solve 11-|2-2x|=x-5 and 11-|2-2x|=5-x and determine whether the solutions are in their correct intervals.

OpenStudy (anonymous):

There are 4 cases, two cases for | x -5| and two cases for | 2 - 2x| : ( x - 5 )+ ( 2 - 2x) = 11 ( x - 5 ) - ( 2 - 2x) = 11 - ( x - 5) + ( 2 - 2x) = 11 - ( x - 5) - ( 2 - 2x) = 11

OpenStudy (anonymous):

@blockcolder Just expand your instruction, hopefully the asker won't be confused with absolute value :)

OpenStudy (anonymous):

You can also consider 3 cases. \[ x \le 1 \\ 1 \le x \le 5\\ \ 5 \le x \]

OpenStudy (anonymous):

For the first case \[ |x-5|=11-|2-2x| \\ -x +5 = 11 -(2-2x)\\ -x+5 = 11 -2 + 2x\\ -3 x =4 \\ x= -\frac 4 3 \] Do the remaining two cases in a similar way/

OpenStudy (anonymous):

Thank you for your replies. I did as you suggest and get four solutions: A. ( x - 5 )+ ( 2 - 2x) = 11, giving x = 14, which does not satisfy the equation. B. ( x - 5 ) - ( 2 - 2x) = 11, giving x = 6, which satisfies the equation. C. - ( x - 5) + ( 2 - 2x) = 11, giving x = -4/3, which satisfies the equation D. - ( x - 5) - ( 2 - 2x) = 11, giving x = 8, which does not satisfy the equation. Am I missing something?

OpenStudy (anonymous):

For the second case \[ |x-5|=11-|2-2x| \\ -x +5 = 11 -(-2+2x)\\ -x+5 = 11+2 -2x\\ x =8 \\ \] But 8 is not in the interval [1,5] so this solution is to be rejected.

OpenStudy (anonymous):

For the third case \[|x-5|=11-|2-2x| \\ x -5 = 11 -(-2+2x)\\ x-5 = 11+2 -2x\\ 3x =18 \\ x = 6 \] So you are right. But, you only needed to consider three cases and not 4

OpenStudy (anonymous):

Thanks, but am still slightly confused. How did you derive the three cases x<1, x between 1 and 5, and x >5 in your reply above? Also, what about the case x = 14? Both 8 and 14 are greater than 5. Why do you exclude them on the grounds that 8 and 14 are not in the interval [1,5]?

OpenStudy (anonymous):

you choose 1, because (2 -2x) changes sign at at x =1. you choose 5, because (x-5) changes sign at at x =5. Since |u|= u if u >=0 and |u| =-u if u<=0.

OpenStudy (anonymous):

thank you so much.

OpenStudy (anonymous):

yw